OpenStudy (anonymous):

y varies directly as x and y=20 when x=3

OpenStudy (anonymous):

y=kx 20 = 3k k=20/3 y = (20/3 ) x

OpenStudy (anonymous):

y=kx 20 = 3k k=20/3 y = (20/3 ) x

OpenStudy (anonymous):

thanks